Solve.
y = -5x – 6
y = -x + 6

Answer by ewatrrr(24785)

 
Hi,
y = -5x – 6
and
y = -x + 6
then ***
-5x-6 = -x + 6 |Adding 5x and -6 to both sides of the EQ
-12 = 4x
-3 = x
CHECKING our answer***
Using -5x-6 = -x + 6
15 - 6 = 3 + 6 = 9